OpenStudy (anonymous):

let p be the midpoint of AB. Find the coordinates of A, Given the coordinates of P and B 1. B(-1,3) P(2,2)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ a_x}}\quad ,&{\color{blue}{ a_y}})\quad B&({\color{red}{ -1}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad \left(\cfrac{{\color{red}{ -1}} + {\color{red}{ a_x}}}{2}\quad ,\quad \cfrac{{\color{blue}{ 3}} + {\color{blue}{ a_y}}}{2} \right)=(2,2)\ P \\ \quad \\ thus\to \begin{cases} \cfrac{{\color{red}{ -1}} + {\color{red}{ a_x}}}{2}-2 \\ \quad \\ \bf \cfrac{{\color{blue}{ 3}} + {\color{blue}{ a_y}}}{2}=2 \end{cases} \)

OpenStudy (jdoe0001):

hmmm shoot a minus typos anyhow if point P is (2,2) that means \(\bf thus\to \begin{cases} \cfrac{{\color{red}{ -1}} + {\color{red}{ a_x}}}{2}=2 \\ \quad \\ \bf \cfrac{{\color{blue}{ 3}} + {\color{blue}{ a_y}}}{2}=2 \end{cases} \)

OpenStudy (jdoe0001):

and you can just solve for \(\bf a_x\ and \ a_y\) to find the coordinates of A

OpenStudy (anonymous):

(-1+a)/2 =2?? i don't get it

OpenStudy (jdoe0001):

well... lemme call those 2 guys ... a different name then.. "a" and b" so we can find the coordinates (a,b) for point A so \(\large thus\to \begin{cases} \cfrac{{\color{red}{ -1}} + {\color{red}{ a}}}{2}=2\to &{\color{red}{ a}}=\square ? \\ \quad \\ \cfrac{{\color{blue}{ 3}} + {\color{blue}{ b}}}{2}=2\to &{\color{blue}{ b}}=\square ? \end{cases} \)

OpenStudy (anonymous):

ooooooh, xD just me being an idiot. xD a=5 b=1

OpenStudy (jdoe0001):

yeap thus A = (5,1)